OpenStudy (anonymous):

the derivative of y=e^3 ln x

OpenStudy (nowhereman):

If you mean \[y=e^{3\ln x}\] use power rules to eliminate e and ln

OpenStudy (anonymous):

no it's \[(e^3)(lnx)\]

OpenStudy (amistre64):

e^3 is jsut a constant so like any constant put it aside and derive ln(x) D(ln(x)) = 1/x now bring the constant back to it... e^3 --- x

OpenStudy (anonymous):

oh, just as if the number 2 was in place of e^3? I would just leave it? oh, okay thanks

OpenStudy (amistre64):

yep..... that e may look like a variable, but its just like "pi" in the sense thatits stands for an actual number :)

OpenStudy (anonymous):

but how would you know if it's a constant? Because originally I used the product rule.

OpenStudy (amistre64):

There is this guy in mathmatics calle Euler. and he has a special number that pops up alot in natural stuff.. 2.71828182845905....... or something like that. So whenever you see an "e" being used in an equation, they are representing Eulers number.

OpenStudy (anonymous):

okay, well thank you

OpenStudy (nowhereman):

Well, it pops-up a lot in mathematics too ^^. \[f(x) = e^x\] is the solution to the initial-value problem \[f'(x) = f(x)\] and also \[e^x = \lim_{n→∞}(1+\frac x n)^n = \sum_{n=0}^∞ \frac{x^n}{n!}\] expanding it to complex numbers you get \[e^{iπ} + 1 = 0\] So it really is a pretty amasing number.

OpenStudy (amistre64):

i use it on my taxes :)

OpenStudy (amistre64):

a better definition for "e" might be: lim(n->inf) (1 + 1/n)^n right?

OpenStudy (amistre64):

whcih of course is whats up there lol..... gotta quit glossing over stuff

OpenStudy (nowhereman):

Hehe, there so many ways to define it :-) Yet another would be to first define \[\ln x = \int_1^x \frac 1 x\] and then say e^x is the inverse function.